The actual cost of direct labour per hour is​ $13.00. Two standard direct labour hours are allowed per finished good. During the current​ period, 2,000 finished goods were produced using​ 3,800 direct labour hours. The direct labour efficiency variance is​ $1,300, favourable. Which one of the following amounts is the standard direct labour rate per hour for the​ period?.

A.

​$6.50

B.

​$10.00

C.

​$13.00

D.

​$26.00
Answer #1

Answer : A = $ 6.50

>> Standard Hours = 2 * 2000 = 4000 hours.

>> Labor Efficiency Variance = ( Standard Hours - Actual Hours ) * Standard Price.

>> $ 1300 = ( 4000 - 3800 ) * Standard Price.

>> Standard Price = $ 6.5